OpenStudy (anonymous):

how do you convert 320km/h to m/s

OpenStudy (anonymous):

 then you should have your answer... remember that when units are multiplied by the reciprocal they cancel

OpenStudy (ahsome):

Simply, divide the km/h by 3.6 320/3.6=88.8m/s

OpenStudy (zehanz):

Or do it like this: 320 km/h is like a division: it is 320 km / 1 h. 320 km = 320000m 1 h = 3600 s, So the division becomes: 320000 m / 3600 s = 320000/3600 m/s. Now carry out that division!

OpenStudy (ahsome):

\[1 KM/H = \frac{1}{3.6}M/S\]

OpenStudy (anonymous):

pretty much what @ZeHanz said... and he said it better than me

OpenStudy (ahsome):

\[1M/S=3.6KM/H\]

OpenStudy (ahsome):

To convert KM/H to M/S, divide by 3.6 To convert M/S to KM/H, multiply by 3.6

OpenStudy (zehanz):

@Ahsome : No matter how BIG you write it: without explanation it does not really help...

OpenStudy (anonymous):

lol

OpenStudy (ahsome):

@ZeHanz Thats how we were taught in Physics. Its just a number you are meant to learn. Like how you know that \(\pi\) is 3.16 etc, you are not meant to know how to get that number in most cases.

OpenStudy (ahsome):

Of course, in some tests it might be a requirement.

OpenStudy (anonymous):

when i learned physics i had to convert the way i showed it at the top

OpenStudy (ahsome):

What I learnt in physics is to remember 3.6 ;)

OpenStudy (anonymous):

we all learn different ways

OpenStudy (ahsome):

Exactly. I was just showing an easier way

OpenStudy (zehanz):

How can you learn physics if you are only taught that some calculation has to be done in a fixed way? Physics or maths, if something can be explained, it should be imo... Once you've seen why the conversion involves a factor 3.6, it is of course handy to use!

OpenStudy (ahsome):

Do you understand, @angela?

OpenStudy (ahsome):

I was shown how to do the conversion. Its just that in practice we use 3.6. We don't do the whole equation.

OpenStudy (ahsome):

You already showed the logic behind it. I was just showing what it meant.

OpenStudy (anonymous):

for someone beginning to learn in that area... whether it be physics or a certain math, it is much better to show work instead of just simplifying it from the start

OpenStudy (ahsome):

I know that @elementwielder. But I also saw you already showed him how to do the work. I just showed the simplified version, knowing you already shown him the expanded version. Thats all :)

OpenStudy (ahsome):

I am not saying you are wrong. I just simplified it.

OpenStudy (zehanz):

I think we all agree! Let's wait to see what @angela10 has to say about it :)

OpenStudy (anonymous):

alright lets not argue anymore and just leave the question be

OpenStudy (ahsome):

He is not online :(

OpenStudy (ahsome):

So we couldn't even help him

OpenStudy (anonymous):

he was online when i responded

OpenStudy (ahsome):

I know. I mean now

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

oh well... hopefully he or she looks at it later and responds

OpenStudy (ahsome):

Yah.

OpenStudy (ahsome):

Not to hijack the thread, but can you guys help me with my problem? We are asked to do Directrix and Focus for our assignment, but our teacher didn't teach it to us yet.

OpenStudy (anonymous):

???????????????????????????????????????

OpenStudy (ahsome):

Sorry, @nawalseif? What's so confusing?